OpenStudy (anonymous):

simplify (((a2)*(8 - 9a2)) /((9)*(4a4 - 10a3 + 9a2 - 24)))

OpenStudy (anonymous):

add all like terms 9a2+9a2= __a2

OpenStudy (anonymous):

i guess it'd be better if i provieded what it actually looks like.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

@scobacat4u

OpenStudy (triciaal):

to divide by a fraction invert the divisor and multiply factor the quadratics and rewrite the problem using the factored form cancel the common factors to simplify

OpenStudy (triciaal):

9-a^2 = difference of 2 squares (3- a)(3 + a) a^2 -6 a + 9 = (a -3)(a - 3) 4 a^2 - 4a - 24 = 4 ( a^2 - a - 6) = 4 (a-3)(a +2) 8 = 4*2

OpenStudy (triciaal):